OpenStudy (anonymous):

2/x + 2/x+1 - 2/x+2 What are the restrictions when simplified? And what's the answer when simplified? Thanks

OpenStudy (saifoo.khan):

\[\frac{2}{x}+\frac{2}{x+1}-\frac{1}{x+2}\]

OpenStudy (saifoo.khan):

is it like that?

OpenStudy (anonymous):

Yes expect the last fraction has a 2 on top :)

OpenStudy (anonymous):

- 2 / x +2

OpenStudy (saifoo.khan):

which 2?

OpenStudy (anonymous):

The last fraction where you put -1 it's -2

OpenStudy (saifoo.khan):

O Alright.

OpenStudy (saifoo.khan):

\[\frac{2 (x^2+4 x+2)}{x (x+1) (x+2)}\]